OpenStudy (mathrocks2020):

Need Someone Gr8 at Math

OpenStudy (mathrocks2020):

The dot plots below show the scores for a group of students for two rounds of a quiz:

OpenStudy (mathrocks2020):

Which of the following inferences can be made using the dot plots? There is no overlap between the data. The range of each round is the same. Round 1 scores are higher than round 2 scores. Round 2 scores are lower than round 1 scores.

OpenStudy (faiqraees):

2

OpenStudy (photon336):

I believe this is showing the frequency # number of instances each score appears.

OpenStudy (mathrocks2020):

What do you mean?

OpenStudy (photon336):

so it's how many times a score appears in each round

OpenStudy (mathrocks2020):

But how do o fond my answer

OpenStudy (mathrocks2020):

@Photon336

OpenStudy (photon336):

so for round 1 how many of those dots are under a score of 4?

OpenStudy (mathrocks2020):

6

OpenStudy (mathrocks2020):

Am I correct?

OpenStudy (mathrocks2020):

He Looo?

OpenStudy (photon336):

yep now each one is four so there are 24 points in 4. for round 1 what about 5 in round 1 too?

OpenStudy (mathrocks2020):

2

OpenStudy (photon336):

Once you find the number of points in round 1 and 2, you can easily eliminate two choices.

OpenStudy (photon336):

so @mathrocks2020 so what are the total number of points for round 1

OpenStudy (mathrocks2020):

8

OpenStudy (photon336):

Round #1 6*4 = 24+ (5*2) = 24+10 = 34

OpenStudy (mathrocks2020):

Still confused on how u got 6*4

OpenStudy (mathrocks2020):

@Photon336 still don't understand how u got 6*4

OpenStudy (photon336):

you see that there are six dots right?

OpenStudy (mathrocks2020):

Oh did u multiply it with the collum it is in?

OpenStudy (photon336):

yep.

OpenStudy (mathrocks2020):

Silly me

OpenStudy (photon336):

so the score for each one of those is 4

OpenStudy (mathrocks2020):

Ok

OpenStudy (photon336):

now i went over to 5 and did the same thing. there are two dots each with a score of 5 so i did 5*2 = 10 now to get the total number of points for round one I did 24+10 = 34

OpenStudy (mathrocks2020):

I see

OpenStudy (mathrocks2020):

Should I try round 2?

OpenStudy (photon336):

now try finding the score for the next round for me

OpenStudy (mathrocks2020):

Round 2:4*4=16+(5*4)=20 16+20=36?

OpenStudy (photon336):

alright great so now. Round 2 = 36 Round 1 = 34 knowing this, now what two choices can you eliminate? let's go to the choices: There is no overlap between the data. The range of each round is the same. Round 1 scores are higher than round 2 scores. Round 2 scores are lower than round 1 scores

OpenStudy (mathrocks2020):

@Photon336 pls don't go:(

OpenStudy (photon336):

i'm still here

OpenStudy (photon336):

@mathrocks2020 are the scores from Round 2 scores lower than round 1 scores? Round 2 36 Round 1 34

OpenStudy (mathrocks2020):

No

OpenStudy (photon336):

we can eliminate that answer choice

OpenStudy (photon336):

What about this: Round 1 scores are higher than round 2 scores.

OpenStudy (mathrocks2020):

Yes then can be eliminated

OpenStudy (photon336):

yep great

OpenStudy (photon336):

last two choices we have are these: There is no overlap between the data. The range of each round is the same.

OpenStudy (photon336):

now do you know what the range is?

OpenStudy (mathrocks2020):

No

OpenStudy (photon336):

the range is the highest value in our data set minus the lowest number

OpenStudy (photon336):

for round 1 the highest number is 5 and the lowest is 4 so 5-4 = 1 what about round 2?

OpenStudy (mathrocks2020):

5-4=1

OpenStudy (photon336):

yep so we know that's true.

OpenStudy (mathrocks2020):

Yes

OpenStudy (photon336):

now the other choice "There is no overlap between the data". is false because there is overlap, the number of data points in both sets are the same so there is overlap.

OpenStudy (mathrocks2020):

Indeed

OpenStudy (mathrocks2020):

So the answer is the range of each data is the same?

OpenStudy (photon336):

yeah it should be the one that makes the most sense.

OpenStudy (mathrocks2020):

Thanks